hi please help me with this problem: The sum of two numers is -17. If one number is one less than three times the other, find the numbers.
Answer by schrammbledeggs(41)

x = number 1
y = number 2

x + y = -17 (equation from statement 1 of problem)
x = 3y - 1 (equation from statement 2 of problem)

Let's plug that value of x into the first equation...
(3y - 1) + y = -17
4y - 1 = -17
4y - 1 + 1 = -17 + 1
4y = -16
y = -16/4
y = -4

Now that we have y we can find x...
Plug that value of y into 1 of those equations (it doesn't matter which one)
x + -4 = -17
x + -4 + 4 = -17 + 4
x = -13

So your numbers are -4 and -13 :)